Find all real numbers that are not in the domain of the following function:
f(x)=x^2+4x+1/2x^2+x-3

Answer by RAY100(1637)

f(x) = (x^2 +4x +1) / (2x^2 +x-3)
.
factor
f(x) = (x+3.732)(x+.268) / (2x+3) (x-1)
.
ps. used quadratic formula for numerator, but it doesn't come into play
.
To be out of the Domain, the function can be (1) undefined, by having zero in denominator,,or(2) complex,,sqrt of negative number
in this case if (2x+3) =0,, x=-3/2,,,,or (x-1)=0,,x=1,,,, function is undefined
.
Therefore Domain is ALL REAL NUMBERS,,except x=-3/2 and x=1
.